# Can Code-Switched Texts Activate a Knowledge Switch in LLMs? A Case Study on English-Korean Code-Switching

### v1

@@ B.2 Prompts
B.3 Human Evaluation Criteria
We provide evaluation guideline for human evaluation.
Faithfulness.
Faithfulness evaluates the factual correctness of the knowledge.
Helpfulness.
Pair-wise comparison.
We comprehensively evaluate the quality of knowledge generated from CS and English question in terms of both faithfulness and helpfulness. If both are identical, evaluators can choose Tie.
Appendix C Additional Results
In this section, we provide additional results and and comprehensive observations throughout our work.
C.1 Knowledge Leveraging Results
| Llama3 8B | CS | 74.58 | 37.50 | 39.15 | 30.13 | 32.41 | 82.14 | 75.61 | 50.50 | 63.06 | 53.90 |
C.2 Error Analysis
| GPT-4o | CS | 0/0 | 0/0 | 0/0 | 0/0 | 2/0 | 0/0 | 0/0 | 5/0 | 1/0 | 8/0 |
| EN | 8/0 | 1/0 | 11/1 | 15/0 | 15/2 | 3/1 | 0/0 | 40/1 | 6/0 | 99/5 | |
| GPT-3.5 | CS | 0/0 | 1/1 | 1/1 | 2/2 | 0/0 | 0/0 | 0/0 | 1/1 | 0/0 | 5/5 |
| EN | 0/0 | 3/3 | 1/1 | 2/2 | 1/1 | 0/0 | 0/0 | 6/6 | 0/0 | 13/13 | |
| Claude 3.5 Sonnet | CS | 0/0 | 0/0 | 0/0 | 0/0 | 0/0 | 0/0 | 0/0 | 0/0 | 0/0 | 0/0 |
| EN | 0/0 | 0/0 | 0/0 | 0/0 | 1/0 | 0/0 | 0/0 | 4/0 | 0/0 | 5/0 | |
| Solar | CS | 0/0 | 5/5 | 1/1 | 2/2 | 11/11 | 0/0 | 0/0 | 5/5 | 1/1 | 20/20 |
| EN | 0/0 | 9/9 | 8/8 | 11/11 | 10/10 | 0/0 | 1/1 | 4/4 | 2/2 | 35/35 | |
| Llama3 70B | CS | 2/2 | 6/6 | 6/6 | 11/11 | 3/3 | 0/0 | 1/1 | 3/3 | 2/2 | 34/34 |
| EN | 2/2 | 7/7 | 10/10 | 24/24 | 12/12 | 2/2 | 0/0 | 4/4 | 2/2 | 63/63 | |
| Llama3 8B | CS | 1/1 | 5/5 | 3/3 | 9/9 | 2/2 | 1/1 | 1/1 | 0/0 | 1/1 | 23/23 |
| EN | 0/0 | 4/4 | 6/6 | 8/8 | 5/5 | 2/2 | 0/0 | 2/2 | 2/2 | 29/29 | |
| Gemma2 27B | CS | 4/4 | 18/18 | 7/7 | 28/26 | 38/22 | 2/2 | 3/2 | 9/6 | 14/10 | 123/97 |
| EN | 3/3 | 28/28 | 5/5 | 38/37 | 15/7 | 7/7 | 1/1 | 20/13 | 5/5 | 122/106 | |
| Gemma2 9B | CS | 3/3 | 7/7 | 12/12 | 25/25 | 19/18 | 1/1 | 2/2 | 7/7 | 2/2 | 78/77 |
| EN | 9/9 | 30/30 | 13/13 | 35/35 | 11/11 | 5/5 | 0/0 | 28/23 | 9/9 | 140/135 |
C.3 Case Study
In this section, we provide a case study of Gemma2 9B on Law domain. In Table 11, hallucinations are observed in the knowledge generated from CS question. According to the Civil Act of the Republic of Korea, individuals under the age of 14 can only enter into binding contracts with the consent of their legal guardians. Additionally, individuals between the ages of 14 and 19 are not deprived of contractual effect; rather, they are granted the right to cancel such agreements at their discretion. Moreover, the knowledge generated in English incorrectly applies the U.S. standard, which defines minors as those under 18 years of age, instead of the Korean standard, which applies to individuals under 19 years of age. This finding suggests that English question is not helpful for identifying necessary and langugae-specific knowledge.
| Example of Solar from the Tradition domain |
| [QUESTION] |
| The following is a question about 한국의 문화 and 전통. |
| (The following is a question about Korean culture and traditions. |
| [Knowledge from CS] |
| [ANSWER] |
| [Knowledge from EN] |
| [ANSWER] |
| (4) … |
| Example of Gemma2 9B from the Law domain |
| [QUESTION] |
| What is the correct legal 판단 on the following case? |
| (What is the correct legal judgment on the following case? |
| (1) B’s 부모 can cancel the 계약. (B’s parents can cancel the contract.) |
| [Knowledge from CS] |
| 2. **부모의 동의**: 미성년자의 계약은 부모의 동의가 있어야 효력이 있다. 동의서가 위조된 경우, 계약은 무효이다. |
| 3. **계약의 취소**: 계약의 당사자는 계약을 취소할 수 있다. |
| 4. **의사표시의 취소**: 의사표시는 당사자의 의사에 따라 취소될 수 있다. |
| [ANSWER] (Correct answer is (3).) |
| [Knowledge from EN] |
| [ANSWER] |

### v2

@@ B.2 Prompts
B.3 Comparing CS and original Korean
We also present experimental results on original Korean data in Table 9. Generally, performances in original Korean are higher than in CS, while CS score approximates or equal to in many cases. Considering the English dominance of the LLMs, GPT-4o and Claude 3.5 Sonnet present advanced multilingual ability with over 80% accuracy. As our main research focus was towards on examining the effect of code-switching compared to dominant language, we exclude original Korean as our major concern. Also, there is a severe possibility that the performances may be influenced by already having seen the datasets, i.e., data contamination, as the original datasets are sourced from official exams and texts from the Internet that are openly available.
B.4 Open-ended QA.
Out dataset, EnKoQA is multiple-choice QA dataset, following its original source datasets. We additionally explore the potential of code-switching on open-ended QA as well.
Results are shown in Table 5. Using same questions in our dataset, we instruct the model to respond in short answer and compute exact match score. It is noticeable that the performances are very low compared to multiple-choice QA results. We attribute this to the free-form response of open-ended tasks, causing more errors and hallucinations. It is observable that the models barely answer correctly in History and Popular.
| Model | Economy | Geography | History | Law | Politics | Popular | Society | Tradition | |
| GPT-4o | CS | 85.00 | 20.00 | 40.00 | 30.00 | 30.00 | 05.00 | 50.00 | 35.00 |
| EN | 80.00 | 00.00 | 05.00 | 05.00 | 10.00 | 00.00 | 05.00 | 00.00 | |
| KOR | 85.00 | 65.00 | 65.00 | 40.00 | 75.00 | 45.00 | 85.00 | 95.00 | |
| GPT-3.5 | CS | 70.00 | 00.00 | 00.00 | 20.00 | 10.00 | 05.00 | 10.00 | 10.00 |
| EN | 75.00 | 00.00 | 00.00 | 10.00 | 15.00 | 0.00 | 05.00 | 00.00 | |
| KOR | 65.00 | 45.00 | 05.00 | 30.00 | 60.00 | 20.00 | 65.00 | 60.00 | |
| Llama3-70B | CS | 20.00 | 00.00 | 00.00 | 10.00 | 15.00 | 10.00 | 10.00 | 00.00 |
| EN | 30.00 | 05.00 | 00.00 | 10.00 | 20.00 | 05.00 | 10.00 | 00.00 | |
| KOR | 60.00 | 50.00 | 00.00 | 40.00 | 70.00 | 35.00 | 55.00 | 60.00 | |
| Llama3-8B | CS | 20.00 | 00.00 | 00.00 | 05.00 | 25.00 | 00.00 | 05.00 | 00.00 |
| EN | 15.00 | 00.00 | 00.00 | 05.00 | 15.00 | 00.00 | 00.00 | 00.00 | |
| KOR | 25.00 | 30.00 | 05.00 | 05.00 | 50.00 | 05.00 | 10.00 | 20.00 |
Appendix C Evaluation Details
C.1 Evaluation Criteria
We provide evaluation guideline for human evaluation.
Faithfulness.
Faithfulness evaluates the factual correctness of the knowledge.
Helpfulness.
Pair-wise comparison.
We comprehensively evaluate the quality of knowledge generated from CS and English questions in terms of both faithfulness and helpfulness. If both are identical, evaluators can choose Tie.
In case of LLM-as-a-judge evaluation, same criteria and instructions are given as prompts.
C.2 Human Evaluator Qualifications
For knowledge identification evaluation, collecting qualified bilingual evaluators was not easy due to the inherent challenge in code-switching research of necessitating fluent bilinguals as evaluators. Our dataset is composed of questions from Korean proficiency tests for foreigners and the Korean College Scholastic Ability Test. Thus, it is designed at a level that would not be challenging for evaluators whom were born and raised in Korea, received a Korean public education, and graduated prestigious universities. We managed to collect four Korean graduate school students as our evaluators, all of whom are native Korean with sufficient understanding of Korean culture. Also, they possess qualified English exam scores, indicating that they have no problem in understanding Korean-English code-switched texts. To mitigate the shortage of labor force, we designed the evaluation criteria objectively, allowing for an assessment that is not subjective and has clear correct answers. Specifically, we evaluate knowledge identification based on two criteria: faithfulness and helpfulness. Faithfulness evaluates the factualness of the knowledge, so the evaluators are required to use their background knowledge as well as searching from faithful sources where gold knowledge exists. To evaluate helpfulness, evaluators are given a gold answer to the question and determine whether the knowledge is helpful for finding the answer, using their logical reasoning.
Appendix D Observations
In this section, we provide additional results and comprehensive observations throughout our work.
| Model | History | Tradition | General | Law |
| GPT-4o | 0.41 | 0.64 | 0.62 | 0.62 |
| Solar | 0.26 | -0.09 | 0.38 | 0.02 |
| Gemma2 27B | 0.25 | 0.52 | 0.17 | 0.34 |
| Gemma2 9B | 0.20 | -0.07 | 0.05 | 0.24 |
D.1 Knowledge Identification Results
We observed that the majority of models benefitted from CS questions. Table 1 shows that scores in CS are higher on all models in Politics, and in case of Law, only three models (GPT-3.5, Llama3 70B, and Gemma2 9B) out of eight models performed worse. We can see in Average score, all models except Gemma2 27B performed better on CS.
D.2 Knowledge Leveraging Results
| Llama3 8B | CS | 76.27 | 39.20 | 40.57 | 38.46 | 33.33 | 72.62 | 73.17 | 47.03 | 56.31 | 53.00 |
| EN | 72.88 | 37.50 | 40.57 | 32.91 | 31.72 | 72.62 | 75.61 | 50.33 | 59.01 | 52.57 | |
D.3 Error Analysis
| GPT-4o | CS | 0/0 | 0/0 | 0/0 | 0/0 | 0/2 | 0/0 | 0/0 | 0/5 | 0/1 | 0/8 |
| EN | 0/8 | 0/1 | 1/11 | 0/15 | 2/15 | 1/3 | 0/0 | 1/40 | 0/6 | 5/99 | |
| GPT-3.5 | CS | 0/0 | 1/1 | 1/1 | 2/2 | 0/0 | 0/0 | 0/0 | 1/1 | 0/0 | 5/5 |
| EN | 0/0 | 3/3 | 1/1 | 2/2 | 1/1 | 0/0 | 0/0 | 6/6 | 0/0 | 13/13 | |
| Claude 3.5 Sonnet | CS | 0/0 | 0/0 | 0/0 | 0/0 | 0/0 | 0/0 | 0/0 | 0/0 | 0/0 | 0/0 |
| EN | 0/0 | 0/0 | 0/0 | 0/0 | 0/1 | 0/0 | 0/0 | 0/4 | 0/0 | 0/5 | |
| Solar | CS | 0/0 | 5/5 | 1/1 | 2/2 | 11/11 | 0/0 | 0/0 | 5/5 | 1/1 | 20/20 |
| EN | 0/0 | 9/9 | 8/8 | 11/11 | 10/10 | 0/0 | 1/1 | 4/4 | 2/2 | 35/35 | |
| Llama3 70B | CS | 2/2 | 6/6 | 6/6 | 11/11 | 3/3 | 0/0 | 1/1 | 3/3 | 2/2 | 34/34 |
| EN | 2/2 | 7/7 | 10/10 | 24/24 | 12/12 | 2/2 | 0/0 | 4/4 | 2/2 | 63/63 | |
| Llama3 8B | CS | 1/1 | 5/5 | 3/3 | 9/9 | 2/2 | 1/1 | 1/1 | 0/0 | 1/1 | 23/23 |
| EN | 0/0 | 4/4 | 6/6 | 8/8 | 5/5 | 2/2 | 0/0 | 2/2 | 2/2 | 29/29 | |
| Gemma2 27B | CS | 4/4 | 18/18 | 7/7 | 26/28 | 22/38 | 2/2 | 2/3 | 6/9 | 10/14 | 97/123 |
| EN | 3/3 | 28/28 | 5/5 | 37/38 | 7/15 | 7/7 | 1/1 | 13/20 | 5/5 | 106/122 | |
| Gemma2 9B | CS | 3/3 | 7/7 | 12/12 | 25/25 | 18/19 | 1/1 | 2/2 | 7/7 | 2/2 | 77/78 |
| EN | 9/9 | 30/30 | 13/13 | 35/35 | 11/11 | 5/5 | 0/0 | 23/28 | 9/9 | 135/140 |
D.4 Case Study
We examine a sample case to compare the capability of code-switching and English on knowledge activation. Table 10 shows the knowledge and answer generated by Solar in Tradition. The question asks about “정월대보름”, a Korean traditional holiday that celebrates the first full moon of lunar new year. We observe that CS question preserves unique terms such as “정월대보름” and ”귀밝이술” in Korean; this helps the model to successfully activate faithful knowledge, consequently leading to the correct answer. However, in the case of English, not only are these cultural nuances lost in English question, but the model misunderstood the question to asking about “Dan-o”, another Korean traditional holiday. Solar lacks in knowledge about “정월대보름” in English, or fails to activate encoded knowledge with its English translation.
We also provide a case of CS failing in knowledge activation in Table 11. In the case of Gemma2 9B on Law domain, hallucinations are observed in the knowledge generated from CS question. According to the Civil Act of the Republic of Korea, individuals under the age of 14 can only enter into binding contracts with the consent of their legal guardians. Additionally, individuals between the ages of 14 and 19 are not deprived of contractual effect; rather, they are granted the right to cancel such agreements at their discretion. Moreover, the knowledge generated in English incorrectly applies the U.S. standard, which defines minors as those under 18 years of age, instead of the Korean standard, which applies to individuals under 19 years of age. This finding suggests that English question is not helpful for identifying necessary and language-specific knowledge.
| GPT-4o | CS | 91.53 | 78.41 | 69.04 | 74.79 | 55.86 | 90.48 | 95.12 | 63.70 | 85.14 | 78.23 |
| KOog | 94.92 | 76.70 | 75.09 | 76.50 | 58.62 | 89.29 | 97.56 | 67.00 | 85.59 | 80.14 | |
| KOog | 71.19 | 45.45 | 37.37 | 29.49 | 30.11 | 70.24 | 60.98 | 47.36 | 54.05 | 49.58 | |
| Claude 3.5 Sonnet | CS | 93.22 | 72.16 | 72.95 | 73.08 | 62.53 | 86.90 | 95.12 | 67.66 | 84.23 | 78.65 |
| KOog | 89.83 | 71.59 | 80.78 | 76.28 | 66.67 | 89.29 | 95.12 | 71.45 | 87.39 | 80.93 | |
| KOog | 84.75 | 51.70 | 55.87 | 64.74 | 43.22 | 82.14 | 82.93 | 55.28 | 76.58 | 66.36 | |
| KOog | 86.44 | 52.27 | 53.38 | 51.50 | 41.84 | 77.38 | 82.93 | 59.41 | 68.92 | 63.79 | |
| KOog | 72.88 | 38.07 | 37.37 | 36.75 | 35.40 | 72.62 | 70.73 | 51.32 | 55.86 | 52.33 | |
| KOog | 83.05 | 45.45 | 48.75 | 45.94 | 41.38 | 77.38 | 80.49 | 56.60 | 67.12 | 60.68 | |
| KOog | 77.97 | 41.48 | 45.91 | 38.89 | 40.92 | 80.95 | 80.49 | 53.80 | 66.22 | 58.51 |
| Example of Solar from the Tradition domain |
| [QUESTION] |
| The following is a question about 한국의 문화 and 전통. |
| (The following is a question about Korean culture and traditions. |
| [Knowledge from CS] |
| [ANSWER] |
| [Knowledge from EN] |
| [ANSWER] |
| (4) … |
| Example of Gemma2 9B from the Law domain |
| [QUESTION] |
| What is the correct legal 판단 on the following case? |
| (What is the correct legal judgment on the following case? |
| (1) B’s 부모 can cancel the 계약. (B’s parents can cancel the contract.) |
| [Knowledge from CS] |
| 2. **부모의 동의**: 미성년자의 계약은 부모의 동의가 있어야 효력이 있다. 동의서가 위조된 경우, 계약은 무효이다. |
| 3. **계약의 취소**: 계약의 당사자는 계약을 취소할 수 있다. |
| 4. **의사표시의 취소**: 의사표시는 당사자의 의사에 따라 취소될 수 있다. |
| [ANSWER] (Correct answer is (3).) |
| [Knowledge from EN] |
| [ANSWER] |